Solve the exponential equation.
2x = 1/2
3x = 1/27
2-x = 16
4x = 0.5
8x = 1/4

Found 2 solutions by Alan3354, lynnlo:
Answer by Alan3354(30993)

Solve the exponential equation.
2x = 1/2
If x is the exponent:
Use ^ for exponents.
2^x = 1/2
x = -1 by inspection.
The long way:
x*log(2) = log(1/2)
x = log(1/2)/log(2)
x = -1
=====================
Do the others the same way.
3x = 1/27
2-x = 16
4x = 0.5
8x = 1/4

Answer by lynnlo(4164)

(1)1/4
(2)1/81
(3)-14
(4)0.125
(5)1/32